OpenStudy (anonymous):

Doing polynomials in class...anyone? −(4x)2

OpenStudy (anonymous):

-93.61702127

Vocaloid (vocaloid):

-(4x)^2 = -16x^2

OpenStudy (mathstudent55):

\(-(4x)^2\) The exponent of 2 applies only to the product inside the parentheses. To raise a product to an exponent, raise each factor of the product to the exponent. \(-(4x)^2 =\) \(= -(4^2 \cdot x^2)\) \(=-(16x^2)\) \(=-16x^2\) Do you understand the steps?